OpenStudy (anonymous):

Simplify the expression: (14X+21)/7

OpenStudy (anonymous):

2x+3/1

OpenStudy (anonymous):

or just 2x+3

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

all of it ? or just 1?

OpenStudy (anonymous):

\[\frac{ A+B }{ C }=\frac{ A }{ C }+\frac{ B }{ C }\]

OpenStudy (anonymous):

uhm

OpenStudy (anonymous):

he just divided every number by 7

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

the answer would be 2x+3

OpenStudy (anonymous):

all of that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem.. but the other guy aleady gave it to u lol

OpenStudy (anonymous):

ohh.... lol